An orchard has 1023 apple trees. The number of rows exceeds the number of trees per row by 2.How many trees are there in each row?
Answer by Boreal(15235)

rows=r
trees in row= r-2
r(r-2)=1023
r^2-2r-1023=0
r=(1/2)(2+/- sqrt (4+4092); sqrt 4096=64
r=33; negative root doesn't work.
33 rows, 31 trees in each row=1023 trees.